OpenStudy (anonymous):

split 20 into two nonnegative numbers x,y such that the product of x and y^2 is a maximum

OpenStudy (anonymous):

>>> for i in range(21): x = i y = 20 - i print(x, '*', y, '^2 =', x * y**2) 0 * 20 ^2 = 0 1 * 19 ^2 = 361 2 * 18 ^2 = 648 3 * 17 ^2 = 867 4 * 16 ^2 = 1024 5 * 15 ^2 = 1125 6 * 14 ^2 = 1176 7 * 13 ^2 = 1183 8 * 12 ^2 = 1152 9 * 11 ^2 = 1089 10 * 10 ^2 = 1000 11 * 9 ^2 = 891 12 * 8 ^2 = 768 13 * 7 ^2 = 637 14 * 6 ^2 = 504 15 * 5 ^2 = 375 16 * 4 ^2 = 256 17 * 3 ^2 = 153 18 * 2 ^2 = 72 19 * 1 ^2 = 19 20 * 0 ^2 = 0